In an experiment to determine the specific latent heat of fusion of ice, an ice cube is dropped into water that is contained in a well-insulated calorimeter of negligible specific heat capacity. The following data are available.


  

Mass of ice cube = 25g

Mass of water = 350g

Initial temperature of ice cube = 0 ̊C

Initial temperature of water = 18 ̊C

Final temperature of water = 12 ̊C

Specific heat capacity of water = 4200Jkg−1K−1

14a. Using the data, estimate the specific latent heat of fusion of ice.


1
Expert's answer
2021-03-08T08:18:33-0500
"m_wc_w(T_i-T)-m_{w,i}c_wT=m_iL_f""L_f=\dfrac{m_wc_w(T_w-T)}{m_i}-m_{w,i}c_wT,"

"L_f=\dfrac{0.35\ kg\cdot4200\ \dfrac{J}{kg\cdot \!^{\circ}C}\cdot(18^{\circ}C-12^{\circ}C)-0.025\ kg\cdot4200\ \dfrac{J}{kg\cdot \!^{\circ}C}\cdot12^{\circ}C}{0.025\ kg},"


"L_f=3.02\cdot10^5\ \dfrac{J}{kg}."